# Hulk: A Universal Knowledge Translator for Human-Centric Tasks

### v4

@@ B-C Objective function details
Image caption
CUHK-PEDES [90] 68,126
@@
Total
42 30,187,836$\approx$30M
TABLE XIX: Statistics of Training Datasets in the ablation setting.
@@
2D pose estimation
COCO [72] 149,813
(2 datasets)
AIC [73] 378,352
@@
Pedestrian detection
CrowdHuman [78] 15,000
Skeleton action
NTU60 [87] 40,091
(2 datasets)
NTU120 [114] 63,026
Image caption
CUHK-PEDES [90] 68,126
@@
Total
21 1,693,302$\approx$1.7M
Appendix C Potential Negative Impacts
Appendix D Details of training datasets
D-A Full setting
As shown in Table XVIII, we utilize 42 publically available datasets to form the training set for training our Hulk model, containing 30,187,836 training samples and covering eight different human-centric tasks, including human parsing, 2D pose estimation, attribute recognition, pedestrian detection, skeleton-based action recognition, image caption, 3D pose estimation, and mesh recovery.
D-B Ablation setting
In ablation studies, to demonstrate the effectiveness of Hulk and further explore the potential of a human-centric foundation model, we conduct several ablations on a smaller training set. In Table XIX, the ablation set consists of 21 datasets with about 1.7M samples, covering all eight tasks.
Appendix E Details of dataset-wise configurations
E-A Training
We provide detailed dataset-wise training configurations in Table XX, including total batch size, batch size per GPU, the number of GPUs, and task weights. In order to better unify different tasks and different datasets, the different datasets of certain tasks (i.e., attribute recognition, pedestrian detection, skeleton-based action, image caption, 3D pose, and mesh recovery) are stacked together as an integrated dataset for training, and the datasets of some other tasks (i.e., human parsinga nd 2D pose estimation) are separately trained on different GPUs with diverse task weights.
E-B Fine-tuning
For the fine-tuning process, we carefully tune the learning rate, total batch size, the number of iterations, and drop path rate, and report the best performances. In detail, the ViT-base backbone (i.e., in Table XXI) and ViT-large backbone (i.e., in Table XXII) are tuned, respectively.
TABLE XX: Hulk joint training setup.
@@ E-B Fine-tuning
TABLE XXI: Detailed finetuning configs with ViT-base backbone for human-centric tasks.
@@
TABLE XXII: Detailed finetuning configs with ViT-large backbone for human-centric tasks.

### v5

@@ G-B Ablation setting
Image caption CUHK-PEDES [114] 68,126
@@
[…]
TABLE XXII: Statistics of Training Datasets in the ablation setting.
@@
2D pose estimation COCO [97] 149,813
(2 datasets) AIC [98] 378,352
@@
Pedestrian detection CrowdHuman [102] 15,000
Skeleton action NTU60 [111] 40,091
(2 datasets) NTU120 [132] 63,026
Image caption CUHK-PEDES [114] 68,126
@@
[…]
Appendix H Details of dataset-wise configurations
H-A Training
We provide detailed dataset-wise training configurations in Table XXIV, including total batch size, batch size per GPU, the number of GPUs, and task weights. In order to better unify different tasks and different datasets, the different datasets of certain tasks (i.e., attribute recognition, pedestrian detection, skeleton-based action, image caption, 3D pose, and mesh recovery) are stacked together as an integrated dataset for training, and the datasets of some other tasks (i.e., human parsing and 2D pose estimation) are separately trained on different GPUs with diverse task weights.
H-B Steps to adjust task weight
(1) In Hulk, there are two steps to adjust task weights:
  • Step1: Make sure different human-centric tasks provide similar gradients to task-shared parameters. Since de-tokenizers in Hulk are not shared across all human-centric tasks, we use the last parameter (Layer Norm) in the decoder of Hulk as a probe during Hulk’s training on the small ablation set. The L2-normed gradients of this probe from different tasks should be similar.
  • Step2: Ablate the generated task weights by $0.1\times$, $0.3\times$, $3\times$, $10\times$, etc., for better performance of all eight tasks. We provide the L2-normed gradients of all eight tasks in Table XXIII. Even after adjusting task weights, the L2-normed gradients of all eight tasks are almost in the same order of magnitude.
TABLE XXIII: For eight tasks, we compute the L2-normed gradient using the last parameter in the decoder of Hulk.
Attribute Caption Skeleton 3D Pose&Mesh Detection 2D pose Parsing
0.93 1.08 0.48 3.83 5.01 3.35 0.83
H-C Fine-tuning
For the fine-tuning process, we carefully tune the learning rate, total batch size, the number of iterations, and drop path rate, and report the best performances. In detail, the ViT-base backbone (i.e., in Table XXV) and ViT-large backbone (i.e., in Table XXVI) are tuned, respectively.
TABLE XXIV: Hulk joint training setup.
@@ H-C Fine-tuning
TABLE XXV: Detailed finetuning configs with ViT-base backbone for human-centric tasks.
@@
TABLE XXVI: Detailed finetuning configs with ViT-large backbone for human-centric tasks.
@@
Appendix I Experimental results with complete comparison methods
Due to the limited length of our manuscript, we provide experimental results with complete comparison methods here.
TABLE XXVII: Pedestrian detection evaluation on CrowdHuman [102] with mAP, MR-2 and JI. Following UniHCP [2], we report the direct eval and fine-tune (FT) results. †indicates using a smaller input resolution with a maximum height/width of 1120.
[…]
Specialist DETR [121] ResNet-50 75.9 73.2 74.4
CrowdDet [122] ResNet-50 90.7 41.4 82.3
V2F-Net [123] ResNet-50 91.0 42.3 -
PEDR [124] ResNet-50 91.6 43.7 83.3
DDETR [125] ResNet-50 91.5 43.7 83.1
Sparse-RCNN [126] ResNet-50 91.3 44.8 81.3
Iter-DDETR [47] ResNet-50 92.1 41.5 84.0
Iter-Sparse-RCNN [47] ResNet-50 92.5 42.6 83.3
Iter-DDETR [47] Swin-L 94.1 37.7 87.1
Pretraining PATH [30] ViT-B 90.9 - -
PATH [30] ViT-L 90.8 - -
Generalist UniHCP [2] ViT-B 90.0 46.6 82.2
UniHCP-FT [2] ViT-B 92.5 41.6 85.8
Hulk† ViT-B 90.7 43.8 84.0
Hulk-FT† ViT-B 92.4 40.7 86.0
Hulk† ViT-L 92.2 40.1 85.8
Hulk-FT† ViT-L 93.0 36.5 87.0
[…]
Method Backbone COCO AIC
Specialist HRNet [43] HRNet-W32 74.4 -
HRNet [43] HRNet-W48 75.1 -
TokenPose-L/D24 [46] HRNet-W48 75.8 -
HRFormer-B [37] HRFormer-B 75.6 -
ViTPose-B [127] ViT-B 75.8 -
ViTPose-B‡ [127] ViT-B 77.1 32.0
ViTPose++-B‡ [130] ViT-B 77.0 31.9
ED-Pose [60] Swin-L 75.8 -
ViTPose-L [127] ViT-L 78.3 -
ViTPose-L‡ [127] ViT-L 78.7 34.5
ViTPose++-L‡ [130] ViT-L 78.6 34.3
Pretraining HCMoCo† [29] HRNet-W48 76.9 -
SOLDIER [28] Swin-B 76.6 -
HAP [31] ViT-B 77.0 32.3
PATH [30] ViT-B 76.3 35.0
PATH [30] ViT-L 77.1 36.3
Generalist UniHCP [2] ViT-B 76.1 32.5
UniHCP-FT [2] ViT-B 76.6 33.6
Hulk ViT-B 77.0 34.5
Hulk-FT ViT-B 77.5 35.6
Hulk ViT-L 78.3 36.3
Hulk-FT ViT-L 78.7 37.1
TABLE XXIX: Human parsing evaluation on Human3.6m [96], LIP [62] and CIHP [99] with mIoU. We compare our method with other SOTA methods, including specialists, generalists, and pretrainings, with an input resolution of $\times480$. †indicated the re-implementation.
Method Backbone H3.6M LIP CIHP
Specialist SNT [165] ResNet-101 - 54.73 60.87
PCNet [166] ResNet-101 - 57.03 61.05
SCHP [104] ResNet-101 - 59.36 -
CDGNet [38] ResNet-101 - 60.30 65.56
Pretraining HCMoCo [29] HRNet-18 62.50 - -
HCMoCo† [29] HRNet-48 66.01 - -
SOLDIER [28] Swin-B - 60.50 -
PATH [30] ViT-B 65.00 61.40 66.80
PATH [30] ViT-L 66.20 62.60 67.50
Generalist UniHCP [2] ViT-B 65.90 63.80 68.60
UniHCP-FT [2] ViT-B 65.95 63.86 69.80
Hulk ViT-B 68.08 63.95 70.58
Hulk-FT ViT-B 68.56 63.98 71.26
Hulk ViT-L 69.31 65.86 72.33
Hulk-FT ViT-L 69.89 66.02 72.68
TABLE XXX: Pedestrian attribute recognition evaluation on PA-100K [101] and RAPv2 [100] with mA reported.
Method Backbone PA-100K RAPv2
Specialist SSC [128] ResNet-50 81.87 -
C-Tran [167] ResNet-101 81.53 -
Q2L [168] TResNetL 80.72 -
L2L [58] ViT-B 82.24 -
DAFL [129] ResNet-50 83.54 81.04
Pretraining SOLIDER [28] Swin-B 86.37 -
HAP [31] ViT-B 86.54 82.91
PATH [30] ViT-B 86.90 83.10
PATH [30] ViT-L 90.80 87.40
Generalist UniHCP [2] ViT-B 79.32 77.20
UniHCP-FT [2] ViT-B 86.18 82.34
Hulk ViT-B 82.85 80.90
Hulk-FT ViT-B 87.85 85.26
Hulk ViT-L 84.36 82.85
Hulk-FT ViT-L 88.97 85.86
TABLE XXXI: Skeleton-based Action Recognition evaluation on NTU60-XSub [132] and FineGYM [112] with accuracy reported.
Method Backbone NTU60 GYM
Specialist ST-GCN [48] GCN 81.5 25.2
AS-GCN [169] GCN 86.8 -
AGCN [49] GCN 88.5 -
Shift-GCN [54] GCN 90.7 -
CrosSCLR [170] GCN 86.2 -
MCC [50] GCN 89.7 -
SCC [51] GCN 88.0 -
UNIK [52] GCN 86.8 -
CTR-GCN [53] GCN 92.4 -
MS-G3D [133] GCN 91.5 92.0
IndRNN [171] RNN 81.8 -
MTCNN [172] CNN 81.1 -
HCN [173] CNN 86.5 -
PoseConv3D [134] CNN 93.1 92.4
SELFYNet [135] CNN - 87.7
ST-TR [55] Transformer 89.9 -
DSTA-Net [56] Transformer 91.5 -
STTFormer [57] Transformer 89.9 -
Pretraining MotionBERT [1] DSTformer 93.0 -
SkeletonMAE [136] GIN - 91.8
Generalist Hulk ViT-B 93.8 91.6
Hulk-FT ViT-B 94.0 92.2
Hulk ViT-L 94.1 92.3
Hulk-FT ViT-L 94.3 93.2
TABLE XXXII: Monocular 3D human pose and mesh recovery evaluation on 3DPW [95] and Human3.6M [96] among image-based methods.
Method Backbone 3DPW Human3.6M
[…]
Specialist HMR [174] ResNet-50 - 130.0 76.7 88.0 56.8
GraphCMR [175] ResNet-50 - - 70.2 - 50.1
SPIN [176] ResNet-50 116.4 96.9 59.2 62.5 41.1
I2LMeshNet [177] ResNet-50 - 93.2 57.7 55.7 41.1
PyMAF [178] ResNet-50 110.1 92.8 58.9 57.7 40.5
ROMP [179] ResNet-50 105.6 89.3 53.5 - -
ROMP [179] HRNet-W32 103.1 85.5 53.3 - -
PARE [137] ResNet-50 99.7 82.9 52.3 - -
METRO [39] ResNet-50 - - - 56.5 40.6
METRO [39] HRNet-W64 88.2 77.1 47.9 54.0 36.7
PARE [137] HRNet-W32 88.6 74.5 46.5 - -
MeshGraphormer [180] HRNet-W64 87.7 74.7 45.6 51.2 34.5
ProHMR [138] ResNet-50 - - 59.8 - 41.2
OCHMR [181] ResNet-50 107.1 89.7 58.3 - -
3DCrowdNet [182] ResNet-50 98.3 81.7 51.5 - -
FastMETRO [40] ResNet-50 90.6 77.9 48.3 53.9 37.3
FastMETRO [40] HRNet-W64 84.1 73.5 44.6 52.2 33.7
CLIFF [139] ResNet-50 81.2 69.0 43.0 47.1 32.7
VisDB [140] ResNet-50 85.5 73.5 44.9 51.0 34.5
Pretraining MotionBERT [1] DSTformer 88.1 76.9 47.2 53.8 34.9
HAP [31] ViT-B 90.1 56.0 106.3 - -
Generalist Hulk ViT-B 79.8 67.0 39.9 43.6 31.9
Hulk-FT ViT-B 80.7 68.9 41.3 44.9 32.0
Hulk ViT-L 77.4 66.3 38.5 40.3 28.8
Hulk-FT ViT-L 79.9 68.3 40.6 41.4 30.2